[Brahms-l] Shift report 20040204 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Wed Feb 04 2004 - 23:52:10 EST
------------------------------------------------------------
Date: Wed Feb  4 23:52:09 2004, Shift: 16:00-24:00
Supervisor: EJ, with: 
Log book#15 pages 48 - 48

Shift Summary:
------------------------------------------------------------
1. There was no beam for the entire shift.  Right now there is probelms with the AGS.
2. No data was taken.
3. The detector appears to be working fine.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Feb 4 23:52:25 2004

This archive was generated by hypermail 2.1.8 : Wed Feb 04 2004 - 23:52:46 EST